Sena invests $250,000 in a city of Louisville bond that pays 7 percent interest. Alternatively, she...

  1. Sena invests $250,000 in a city of Louisville bond that pays 7 percent interest. Alternatively, she could have invested the $250,000 in a bond recently issued by Microsoft, Inc. that pays 9 percent interest with similar risk as the city of Louisville bond. Assume that Sena's marginal tax rate is 24 percent.

    What is Sena's after-tax rate of return on the city of Louisville bond?

    5.32 percent

    9.00 percent

    6.84 percent

    7.00 percent

    None of the choices are correct

  2. Aimee has found it very difficult to repay her loans. Because of these difficulties, her creditor decided to forgive one of her most recent loans, an amount of $45,000. After the loan was discharged, Aimee had total assets of $232,000 and her remaining loans totaled $217,000. What amount must Aimee include in her gross income?

    $30,000

    $45,000

    $15,000

    $28,000

    $0—Aimee was not solvent when the loan was discharged

  3. Regardless of when a divorce agreement is executed, alimony is included in gross income of the recipient and is deductible for AGI by the payer.

    True

    False

  4. When an attorney provides $100 of services in exchange for $100 of grass cutting services, the atorney does not have $100 of income because cash was not received.

    True

    False
Answer #1

1. Sena's after-tax rate of return on the city of Louisville bond?

5.32 percent

Computation:-

Interest Income from Investment ($250,000 *7%) = $17,500

Less: Tax @ 24% $4,200

After Tax return $13,300

After-tax rate of return= 13,300/250,000*100 = 5.32%

2. Amount must Aimee include in her gross income:

$45,000

Explanation (Law): In general, if you have cancellation of debt income because your debt is canceled, forgiven, or discharged for less than the amount you must pay, the amount of the canceled debt is taxable and you must report the canceled debt on your tax return for the year the cancellation occurs. The canceled debt isn't taxable, however, if the law specifically allows you to exclude it from gross income.

3. Alimony is included in gross income of the recipient and is deductible for AGI by the payer.

True

4. When an attorney provides $100 of services in exchange for $100 of grass cutting services, the atorney does not have $100 of income because cash was not received.

False

Explanation (Law): Exchanges of services result in taxable income for both parties.